Need help solving for x:
2^(3x-2) = 3^(2x+1)

Answer by stanbon(75887)

Need help solving for x:
2^(3x-2) = 3^(2x+1)
-----
Take the log of both sides:
(3x-2)*log(2) = (2x+1)*log(3)
----
Get the x-terms together:
3x*log(2) - (2x)*log(3) = log(3)-2log(2)
---
x[3log(2) - 2log(3)] = log(3)-2log(2)
---
x = log[3/4]/log[8/9]
---
x = 2.4425
=================
Cheers,
Stan H.
===============